OpenStudy (anonymous):

an aircraft is about to take off from a runway which is 850m long its take off speed must be 90m.s calculate the minimum acceleration necessary for the aircraft to take off

OpenStudy (ash2326):

We are given the distance the aircraft covers before it achieves a speed of 90m/s as 850 m Now initially velocity of the aircraft is u 0 m/s final velocity v= 90m/s distance s= 850 m/s a= acceleration? We'll use the relation \[v^2-u^=2as\] u=0 \[(90)^2=2 \times a \times 850\] Can you find acceleration now?

OpenStudy (ash2326):

Sorry it's \[v^2-u^2=2as\]

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

ya makes sense then you square root the answere right

OpenStudy (ash2326):

No we need to find a!!! this does not require square rooting \[2as=v^2\] \[a=\frac{v^2}{2s} m/s^2\]

OpenStudy (anonymous):

oh ya think i get it mind just checking my answer

OpenStudy (ash2326):

Yeah, what do you get?

OpenStudy (anonymous):

4.8 m/s

OpenStudy (ash2326):

It's right but the unit of acceleration is m/s^2 so 4.8 m/s^2

OpenStudy (anonymous):

yeah i know just couldn't figure out how to add superscripts but thanks a million man you were really helpful :D

OpenStudy (ash2326):

Welcome:D